2 ∫ 0 (x 2 + 4) dx

20 (x2 + 4) dx

Grade:12

2 Answers

Anusha
9 years ago
Ans:- 20 (x2 +4)dx=[(x2+1 /2+1 )+4x]   --------- since (xn+1 /n+1)
      [(x3 /3)+4x]02 =[(23/3)+8]=[(8/3)+3]=(8+24)/3=32/3
Anusha
9 years ago
Ans:- 20 (x2 +4)dx=[(x2+1 /2+1 )+4x]   --------- since (xn+1 /n+1)
      [(x3 /3)+4x]02 =[(23/3)+8]=[(8/3)+3]=(8+24)/3=32/3